find the slop of a line through points (-1,3) (2,7)
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Use the formula for the slope:



where and are the coordinates of the given points. Just plug in the numbers and do the arithmetic.

John


Answer by stanbon(75887)

find the slope of a line through points (-1,3) (2,7)
------------
slope = (7-3)/(2--1) = 4/3
=============================
Cheers,
Stan H.